OpenStudy (anonymous):

Is it possible to have f'(3)=8 and f''(3)=-3, explain why

OpenStudy (mathmate):

Correction: Let f(x)=ax^2+bx+c f'(x)=2ax+b f"(x)=2a f"(3)=-3 (for any value of x) So 2a=-3 => a=-3/2 f'(3)=2a(3)+b=8 -9+b=8 b=17 Therefore f(x)=-(3/2)x^2+17x+C satisfies the given conditions (C=constant)